OpenStudy (anonymous):

What is the mass of 20.2cm^3 of gold if the density is 19.3 g/cm^3

OpenStudy (anonymous):

\[\frac {19.3g}{1cm^3} \times 20.2cm^3 \] \[\frac {(19.3g ) (20.2cm^3) }{1cm^3} \] cancel the units as appropriate. I don't have a calculator and don't feel like doing the arithmetic.